Use direct proof to prove that the sum of two odd integers say a and b is even. Note
that we have discussed an odd number to be represented by 2n+1 and even as 2n.

Answer by ikleyn(52784)

.
Use direct proof to prove that the sum of two odd integers say a and b is even. Note
that we have discussed an odd number to be represented by 2n+1 and even as 2n.
~~~~~~~~~~~~~~

Let "a" and "b" be odd integer numbers.


Then, as you discussed earlier,  a = 2n+1, b = 2m+1  for some integer numbers "m" and "n".


It implies that  a + b = (2n+1) + (2m+1) = 2n + 1 + 2m + 1 = 2n + 2m + 2 = 2*(n+m+1).


Thus the sum a + b is an integer number of the form 2k, where integer number "k" is n+m+1.


It means that  a + b  is an even number.

The proof is complete.